# SITSMamba for Crop Classification based on Satellite Image Time Series

### v1

@@ Abstract
Satellite image time series (SITS) data provides continuous observations over time, allowing for the tracking of vegetation changes and growth patterns throughout the seasons and years. Numerous deep learning (DL) approaches using SITS for crop classification have emerged recently, with the latest approaches adopting Transformer for SITS classification. However, the quadratic complexity of self-attention in Transformer poses challenges for classifying long time series. While the cutting-edge Mamba architecture has demonstrated strength in various domains, including remote sensing image interpretation, its capacity to learn temporal representations in SITS data remains unexplored. Moreover, the existing SITS classification methods often depend solely on crop labels as supervision signals, which do not sufficiently emphasize the importance of temporal information for the network. In this paper, we proposed a Satellite Image Time Series Mamba (SITSMamba) method for crop classification based on remote sensing time series data. The proposed SITSMamba contains a spatial encoder based on Convolutional Neural Networks (CNN) and a Mamba-based temporal encoder. To exploit richer temporal information from SITS, we design two branches of decoder used for different tasks. The first branch is a crop Classification Branch (CBranch), which includes a ConvBlock to decode the feature to a crop map. The second branch is a SITS Reconstruction Branch that uses a Linear layer to transform the encoded feature to predict the original input values. Furthermore, we design a Positional Weight (PW) applied to the RBranch to help the model learn rich latent knowledge from SITS. We also design two weighting factors to control the balance of the two branches during training. When evaluated on the PASTIS and MTLCC datasets, our SITSMamba achieves an OA of 0.7416 and 0.9104, respectively, outperforming the previous state-of-the-art (SOTA) methods. The code of SITSMamba is available at: https://github.com/XiaoleiQinn/SITSMamba.
@@ I Introduction
To utilize the temporal information for crop type classification, a series of traditional methods including threshold-based approaches and pre-defined mathematical equations have been proposed [6, 7, 8]. Moreover, machine learning models such as Support Vector Machine (SVM), Random Forest (RF), have also been adopted for identifying crop types [9]. Although these machine learning models can extract phenology information of crops, their performances heavily rely on manual feature selection and are difficult to apply to other regions.
@@
State-space models (SSMs) [18, 19] have recently become popular due to their capability to efficiently handle continuous long-sequence data analysis, which is a common challenge in various fields such as natural language processing and computer vision. As a typical type of SSM, the S4 model has been further advanced with the introduction of Mamba [20], which incorporates a selective mechanism allowing the model to select relevant information based on the input. This design enhances its computational efficiency compared to the widely used Transformer model when dealing with long sequence data. Mamba’s efficiency of processing long sequences makes it a promising candidate for applications in SITS analysis. However, to the best of our knowledge, no study has yet explored the application of Mamba in the classification of SITS data.
The complex temporal pattern of the crop makes it difficult for models to extract robust temporal features, which can easily lead to model overfitting and poor generalization. If the phenological features of two crop categories are similar, misclassification is likely to occur. Although deep learning models are continuously improving and achieving higher accuracy in remote sensing image time series classification tasks, there are still instances of misclassification across categories. This challenge often stems from the similar phenological patterns across various crop types. Moreover, in the previous fully-supervised methods, only category labels are used to supervise the spatiotemporal feature learning (as shown in Fig. LABEL:introframework(a). However, it may be difficult to map a long and complex time series to a category. A natural question arises: apart from crop label, is there any other information can be used for supervising the model to learn robust representation from the complex temporal patterns?
@@ II-B2 Application of Mamba in Remote Sensing
Recently, the Mamba architecture has been successfully integrated into remote sensing applications. Specifically, Chen et al. [33] adapted Mamba for remote sensing change detection tasks through three specialized frameworks. These frameworks surpass current CNN and Transformer approaches on benchmark datasets, showcasing the Mamba architecture’s effectiveness in capturing spatio-temporal relationships for various change detection scenarios. Zhao et al. [34] proposed a Remote Sensing Mamba (RSM) model for dense prediction tasks in large very-high-resolution (VHR) remote sensing images, featuring an omnidirectional selective scan module (OSSM) for capturing global context. Apart from VHR data, Mamba structure has also been applied to hyper-spectral image (HSI) [35, 36]. [37] proposed a MambaHSI model to enhance hyper-spectral image classification by designing specialized spatial and spectral Mamba blocks for integrated feature extraction. This approach demonstrates superior performance and efficiency in HSI classification tasks across various datasets. As for low-level tasks, [38] proposed SSUMamba, a novel denoising model for hyperspectral images that utilizes the spatial-spectral continuous scan (SSCS) Mamba architecture to effectively model long-range dependencies and correlations.
@@ III-A Overview Architecture
An overview of the proposed SITSMamba is shown in Fig. 2. The framework contains three main components: embedding layer, encoder backbone, and segmentation head. The proposed SITSMamba adopts CNN as the spatial encoder and Mamba block as the temporal encoder. The spatial encoder is a Convolutional block (ConvBlock), which is used in the UTAE model [16]. This ConvBlock extracts spatial feature simultaneously from multi-temporal images. It contains a series of convolutional layers, Rectified Linear Unit (ReLU) activations, and normalization layers. Then the SITS data is reshaped and input to the Mamba block. The framework includes two branches, a Classification Branch (CBranch) and a Reconstruction Branch (RBranch). The CBranch uses a max pooling on temporal dimension and generates a mono-temporal image after reshaping. This step aims to produce a feature map with rich information related to crop types. To enhance the spatial and temporal encoding, we design an RBranch which aims to transform the encoded feature map to the original SITS. Both the Classification Branch and the RBranch are used in the training process, whereas during inference, only the CBranch is used for crop classification and the RBranch is ignored.
@@ III-B SSM mechanism
where $\mathrm{\mathbf{A}}\in\mathcal{R}^{N\times N}$, $\mathrm{\mathbf{B}}\in\mathcal{R}^{N\times 1},\mathrm{\mathbf{C}}\in\mathcal{R}^{1\times N}$ are the system matrices. Specifically, the matrix $\mathbf{A}$ dictates the progression of the state vector h(t). The projection matrix $\mathbf{B}$ specifies how much the input $x\left(t\right)$ influences the hidden state. Through the projection parameter matrix $\mathbf{C}$, the hidden state is converted into the output.
S4 model incorporate a timescale parameter, denoted as $\Delta$, to discretize the continuous system parameters $\mathbf{A}$ and $\mathbf{B}$ into their discrete equivalents $\overline{\mathrm{\mathbf{A}}}$ and $\overline{\mathrm{\mathbf{B}}}$. This process aims to adapt the continuous system into DL methods. This transformation typically employs the zero-order hold (ZOH) method that can be expressed as follows:
@@ III-F Architecture detail
Serving as the spatial encoder, the ConvBlock consists of two Convolutional (Conv2D) layers, each applying a 3×3 convolutions with a padding of 1. This parameter setting preserves the spatial resolution of the input data. Both the Conv2D layers have 128 filters. Following each convolution is a batch normalization (BN) layer and a rectified linear unit (ReLU) activation function. The Mamba block is configured with its default parameter settings. For an input SITS data represented as $\mathrm{\mathbf{I}}\in\mathcal{R}^{T\times C\times H\times W}$, the ConvBlock applies convolutional calculation to each temporal frame, focusing on extracting the spatial features. The output feature of ConvBlock has the shape $T\times C_{1}\times H\times W$ where $C_{1}$ is 128 after the final convolution. This feature is then reshaped to $(H\times W)\times T\times C_{1}$, representing a set of $H\times W$ time series with length $T$ and $C_{1}$ channels. It is fed into the Mamba block to encode the temporal context and produce an output of the same shape. This output feature embedded with useful spatiotemporal information is then processed by two branches. For CBranch, a Max pooling layer is applied along the temporal dimension, generating a feature with the shape $H\times W\times C_{1}$. This feature is then fed into a ConvBlock, which is composed of a Conv2D layer that has $C_{1}$ filters followed by a BN layer and a ReLU layer. After this ConvBlock, the feature is decoded to generate a crop map with a shape of $H\times W\times C_{cls}$, where $C_{cls}$ is the number of crop categories. For RBranch, we use a Linear layer to decode the feature from Mamba and predict the original SITS. The training and testing procedures are summarized in the Algorithm 1.
@@ IV-A Datasets
1) PASTIS Dataset [16]: Panoptic Agricultural Satellite TIme-Series (PASTIS) dataset is composed of four sub-regions in France. This dataset comprises 2,433 one-square-kilometer patches, spanning across four Sentinel-2 tile areas within France. Each patch is equipped with a series of multi-spectral satellite images, each featuring 10 bands and measured at a 10 m resolution of 128 × 128 pixels. This dataset contains 20 categories, which poses a challenge for classification task. The imagery was captured by Sentinel-2 satellites between September 2018 and November 2019, with the number of observations varying from 38 to 61 for each site. Leveraging the spatial and descriptive data of all parcels from the French Land Parcel Identification System (FLPIS), the PASTIS dataset annotates each image with corresponding semantic labels. The benchmark includes five folds, and We adopt fold-1 for training, fold-4 for validation and fold-5 for testing. The number of pixels for each class in the PASTIS dataset is shown in Table I.
TABLE I: Statistics of classes in PASTIS dataset
@@ IV-D Implementation details
PASTIS dataset features larger image size and longer temporal length. Therefore, we clip the image into a size of 32×32 pixels as the input for the models. Additionally, padding is adopted to ensure that the input SITS at different locations in a mini-batch have the same temporal length. As for MTLCC dataset, we use the original 48×48 image size as the input. We sample the SITS in the temporal dimension to ensure that each sample in a mini-batch has the same length in the temporal dimension. Given the variable input image sizes across datasets, we adjust the batch size to 8 for the PASTIS dataset, 4 for MTLCC datasets. The Adam optimizer is used to train the model. The learning rate is set to 0.0001. All runs are conducted on NVIDIA GeForce RTX 3090 GPU.
@@ IV-E Result
Figure 4: Qualitative evaluation of the models on PASTIS dataset.
Table III and Table LABEL:mtlccacc show the quantitative comparison results on PASTIS dataset and MTLCC dataset, respectively. As shown in Table III, SITSMamba outperforms other methods in all the three metrics. Among the other methods, UTAE achieves the best result. TSVIT achieves a result slightly lower than UTAE but outperforms the other methods that use LSTM as the temporal encoder. Apart from the overall evaluation, we also show the F1 score on all the crop types in Table  V. SITSMamba obtains the highest accuracy on most classes, especially for those classes hard to classify, such as class 6, 10, 12, and 18.
TABLE III: Quantitative comparison results on PASTIS Dataset.
@@
TABLE V: F1 score of each crop type on PASTIS Dataset.
@@
We present the crop mapping results of the competing methods on the PASTIS dataset in Fig. 4. In areas with complex field shapes and fragmented parcels, SITSMamba has better prediction results than the competing methods. As shown in the red rectangles, SITS can accurately predict the corn, winter rapeseed, grapevine, and mixed cereal. Among the competing methods, BCLSTM can accurately predict the crop types in a parcel, but it fails to delineate the correct boundary. 3D U-Net generates a more fragile result, which is obvious in the meadow parcel in row (b) and the sunflower parcel in row (c). This may be due to its insufficient temporal encoding using CNN. The other LSTM-based methods, BUnetConvLSTM and BAtrousConvLSTM, also generate many tiny wrong parcels. On the contrary, the Transformer-based methods such as UTAE and TSVIT produce better prediction maps. This result is consistent with the quantitative comparison in Table III that the Transformer-based methods surpass the LSTM-based methods.
@@ IV-F Ablation study
Figure 6: F1 scores of crop types with different methods on PASTIS dataset.

### v2

@@ Abstract
Satellite image time series (SITS) data provides continuous observations over time, allowing for the tracking of vegetation changes and growth patterns throughout the seasons and years. Numerous deep learning (DL) approaches using SITS for crop classification have emerged recently, with the latest approaches adopting Transformer for SITS classification. However, the quadratic complexity of self-attention in Transformer poses challenges for classifying long time series. While the cutting-edge Mamba architecture has demonstrated strength in various domains, including remote sensing image interpretation, its capacity to learn temporal representations in SITS data remains unexplored. Moreover, the existing SITS classification methods often depend solely on crop labels as supervision signals, which do not sufficiently emphasize the importance of temporal information for the network. In this paper, we proposed a Satellite Image Time Series Mamba (SITSMamba) method for crop classification based on remote sensing time series data. The proposed SITSMamba contains a spatial encoder based on Convolutional Neural Networks (CNN) and a Mamba-based temporal encoder. To exploit richer temporal information from SITS, we design two branches of decoder used for different tasks. The first branch is a crop Classification Branch (CBranch), which includes a ConvBlock to decode the feature to a crop map. The second branch is a SITS Reconstruction Branch that uses a Linear layer to transform the encoded feature to predict the original input values. Furthermore, we design a Positional Weight (PW) applied to the RBranch to help the model learn rich latent knowledge from SITS. We also design two weighting factors to control the balance of the two branches during training. When evaluated on the PASTIS32 and MTLCC datasets, our SITSMamba achieves an OA of 0.7416 and 0.9104, respectively, outperforming the previous state-of-the-art (SOTA) methods. The code of SITSMamba is available at: https://github.com/XiaoleiQinn/SITSMamba.
@@ I Introduction
To utilize the temporal information for crop type classification, a series of traditional methods including threshold-based approaches and pre-defined mathematical equations have been proposed [6, 7, 8]. Moreover, machine learning models such as Support Vector Machine (SVM), and Random Forest (RF), have also been adopted for identifying crop types [9]. Although these machine learning models can extract phenology information of crops, their performances heavily rely on manual feature selection and are difficult to apply to other regions.
@@
State-space models (SSMs) [18, 19] have recently become popular due to their capability to efficiently handle continuous long-sequence data analysis, which is a common challenge in a wide range of fields, including natural language processing and computer vision. As a typical type of SSM, the S4 model has been further advanced with the introduction of Mamba [20], which incorporates a selective mechanism allowing the model to select relevant information based on the input. This design enhances its computational efficiency compared to the widely used Transformer model when dealing with long sequence data. Mamba’s efficiency of processing long sequences makes it a promising candidate for applications in SITS analysis. However, to the best of our knowledge, no study has yet explored the application of Mamba in the classification of SITS data.
Although deep learning models are continuously improving and achieving higher accuracy in remote sensing image time series classification tasks, there are still instances of misclassification across categories. This challenge often stems from the similar phenological patterns across various crop types [21]. Moreover, the complex temporal pattern of the crop makes it difficult for models to extract robust temporal features, which can easily lead to model overfitting and poor generalization. In the previous fully-supervised methods, only category labels are used to supervise the spatiotemporal feature learning (as shown in Fig. LABEL:introframework(a). However, it may be difficult to map a long and complex time series to a category. A natural question arises: apart from crop label, is there any other information can be used for supervising the model to learn robust representation from the complex temporal patterns?
@@ II-B2 Application of Mamba in Remote Sensing
Recently, the Mamba architecture has been successfully integrated into remote sensing applications. Specifically, Chen et al. [33] adapted Mamba for remote sensing change detection tasks through three specialized frameworks. These frameworks surpass current CNN and Transformer approaches on benchmark datasets, showcasing the Mamba architecture’s effectiveness in capturing spatio-temporal relationships for various change detection scenarios. Zhao et al. [34] proposed a Remote Sensing Mamba (RSM) model for dense prediction tasks in large very-high-resolution (VHR) remote sensing images, featuring an omnidirectional selective scan module (OSSM) for capturing global context. Apart from VHR data, Mamba structure has also been applied to hyperspectral image (HSI) [35, 36]. For example, [37] proposed a MambaHSI model to enhance hyperspectral image classification by designing specialized spatial and spectral Mamba blocks for integrated feature extraction. This approach demonstrates superior performance and efficiency in HSI classification tasks across various datasets. As for low-level tasks, [38] proposed SSUMamba, a novel denoising model for hyperspectral images that utilizes the spatial-spectral continuous scan (SSCS) Mamba architecture to effectively model long-range dependencies and correlations.
@@ III-A Overview Architecture
An overview of the proposed SITSMamba is shown in Fig. 2. The framework contains three main components: embedding layer, encoder backbone, and segmentation head. The proposed SITSMamba adopts CNN as the spatial encoder and Mamba block as the temporal encoder. The spatial encoder is a Convolutional block (ConvBlock), which is used in the UTAE model [16]. This ConvBlock extracts spatial feature simultaneously from multi-temporal images. It contains a series of convolutional layers, Rectified Linear Unit (ReLU) activations, and normalization layers. Then the SITS data is reshaped and input to the Mamba block. The framework includes two branches, a Classification Branch (CBranch) and a Reconstruction Branch (RBranch). The CBranch uses a max pooling on temporal dimension and generates a mono-temporal image after reshaping. This step aims to produce a feature map with rich information related to crop types. To enhance the spatial and temporal encoding, we design an RBranch which aims to transform the encoded feature map to the original SITS. Both the CBranch and the RBranch are used in the training process, whereas during inference, only the CBranch is used for crop classification and the RBranch is ignored.
@@ III-B SSM mechanism
where $\mathrm{\mathbf{A}}\in\mathcal{R}^{N\times N}$, $\mathrm{\mathbf{B}}\in\mathcal{R}^{N\times 1},\mathrm{\mathbf{C}}\in\mathcal{R}^{1\times N}$ are the system matrices. Specifically, the matrix $\mathbf{A}$ represents the progression of the state vector h(t). The projection matrix $\mathbf{B}$ specifies how much the input $x\left(t\right)$ influences the hidden state. Through the projection parameter matrix $\mathbf{C}$, the hidden state is converted into the output.
S4 model incorporates a timescale parameter, denoted as $\Delta$, to discretize the continuous system parameters $\mathbf{A}$ and $\mathbf{B}$ into their discrete equivalents $\overline{\mathrm{\mathbf{A}}}$ and $\overline{\mathrm{\mathbf{B}}}$. This process aims to adapt the continuous system into DL methods. This transformation typically employs the zero-order hold (ZOH) method that can be expressed as follows:
@@ III-F Architecture detail
Serving as the spatial encoder, the ConvBlock consists of two Convolutional (Conv2D) layers, each applying a 3×3 convolutions with a padding of 1. This parameter setting preserves the spatial resolution of the input data. Both the Conv2D layers have 128 filters. Following each convolution is a batch normalization (BN) layer and a rectified linear unit (ReLU) activation function. The Mamba block is configured with its default parameter settings. For an input SITS data represented as $\mathrm{\mathbf{I}}\in\mathcal{R}^{T\times C\times H\times W}$, the ConvBlock applies convolutional calculation to each temporal frame, focusing on extracting the spatial features. The output feature of ConvBlock has the shape $T\times C_{1}\times H\times W$ where $C_{1}$ is 128 after the final convolution. This feature is then reshaped to $(H\times W)\times T\times C_{1}$, representing a set of $H\times W$ time series with length $T$ and $C_{1}$ channels. It is fed into the Mamba block to encode the temporal context and produce an output of the same shape. This output feature embedded with useful spatiotemporal information is then processed by two branches. For CBranch, a Max pooling layer is applied along the temporal dimension, generating a feature with the shape $H\times W\times C_{1}$. This feature is then fed into a ConvBlock, which is composed of a Conv2D layer that has $C_{cls}$ filters followed by a BN layer and a ReLU layer, where $C_{cls}$ is the number of crop categories. After this ConvBlock, the feature is decoded to generate a crop map with a shape of $H\times W\times C_{cls}$. For RBranch, we use a Linear layer to decode the feature from Mamba and predict the original SITS. The training and testing procedures are summarized in the Algorithm 1.
@@ IV-A Datasets
1) PASTIS Dataset [16]: Panoptic Agricultural Satellite TIme-Series (PASTIS) dataset is composed of four sub-regions in France. This dataset comprises 2,433 one-square-kilometer patches, spanning across four Sentinel-2 tile areas within France. Each patch is equipped with a series of multi-spectral satellite images, each featuring 10 bands and measured at a 10 m resolution of 128 × 128 pixels. This dataset contains 20 categories, which poses a challenge for classification task. The imagery was captured by Sentinel-2 satellites between September 2018 and November 2019, with the number of observations varying from 38 to 61 for each site. Leveraging the spatial and descriptive data of all parcels from the French Land Parcel Identification System (FLPIS), the PASTIS dataset annotates each image with corresponding semantic labels. The benchmark includes five folds, and we adopt fold-1 for training, fold-4 for validation and fold-5 for testing. PASTIS dataset features larger image size and longer temporal length. Therefore, we clip the image into a size of 32×32 pixels as the input for the models. We refer to this clipped dataset as PASTIS32. The number of pixels for each class in the PASTIS32 dataset is shown in Table I. All the 20 classes are used for training and evaluation in this study.
TABLE I: Statistics of classes in PASTIS32 dataset
@@ IV-D Implementation details
On the PASTIS32 dataset, padding is adopted to ensure that the input SITS at different locations in a mini-batch have the same temporal length. As for MTLCC dataset, we sample 30 images from the SITS in the temporal dimension to ensure that each sample in a mini-batch has the same length in the temporal dimension. Given the variable input image sizes across datasets, we adjust the batch size to 8 for the PASTIS32 dataset, 4 for MTLCC datasets. All the models are trained for 100 epochs. The Adam optimizer is used to train the model. The learning rate is set to 0.0001. All runs are conducted on NVIDIA GeForce RTX 3090 GPU.
@@ IV-E Result
Figure 4: Qualitative evaluation of the models on PASTIS32 dataset.
Table III and Table LABEL:mtlccacc show the quantitative comparison results on PASTIS32 dataset and MTLCC dataset, respectively. As shown in Table III, SITSMamba outperforms other methods in all the three metrics. Among the other methods, UTAE achieves the best result. TSVIT achieves a result slightly lower than UTAE but outperforms the other methods that use LSTM as the temporal encoder. Apart from the overall evaluation, we also show the F1 score on all the crop types in Table  V. SITSMamba obtains the highest accuracy on most classes, especially for those classes hard to classify, such as class 6, 10, 12, and 18.
TABLE III: Quantitative comparison results on PASTIS32 Dataset.
@@
TABLE V: F1 score of each crop type on PASTIS32 Dataset.
@@
We present the crop mapping results of the competing methods on the PASTIS32 dataset in Fig. 4. In areas with complex field shapes and fragmented parcels, SITSMamba has better prediction results than the competing methods. As shown in the red rectangles, SITS can accurately predict various parcels, including the winter barley, spring barley, beet, and winter durum wheat. Among the competing methods, BCLSTM can accurately predict the crop types of a parcel, but it often fails to delineate the correct boundary. 3D U-Net generates a more fragile result, which is obvious in the soft winter wheat in row (c) and the grapevine parcel in row (d). This may be due to its insufficient temporal encoding using CNN. The other LSTM-based methods, BUnetConvLSTM and BAtrousConvLSTM, also generate many tiny wrong parcels. On the contrary, the Transformer-based methods, that is, UTAE and TSVIT, produce better prediction maps. This result is consistent with the quantitative comparison in Table III that the Transformer-based methods surpass the LSTM-based methods.
@@ IV-F Ablation study
Figure 6: F1 scores of crop types with different methods on PASTIS32 dataset.